# Translating Images to Road Network: A Sequence-to-Sequence Perspective

### v2

@@ Abstract
The extraction of road network is essential for the generation of high-definition maps since it enables the precise localization of road landmarks and their interconnections. However, generating road network poses a significant challenge due to the conflicting underlying combination of Euclidean (e.g., road landmarks location) and non-Euclidean (e.g., road topological connectivity) structures. Existing methods struggle to merge the two types of data domains effectively, but few of them address it properly. Instead, our work establishes a unified representation of both types of data domain by projecting both Euclidean and non-Euclidean data into an integer series called RoadNet Sequence. Further than modeling an auto-regressive sequence-to-sequence Transformer model to understand RoadNet Sequence, we decouple the dependency of RoadNet Sequence into a mixture of auto-regressive and non-autoregressive dependency. Building on this, our proposed non-autoregressive sequence-to-sequence approach leverages non-autoregressive dependencies while fixing the gap towards auto-regressive dependencies, resulting in success on both efficiency and accuracy. We further identify two main bottlenecks in the current RoadNetTransformer on a non-overfitting split of the dataset: poor landmark detection limited by the BEV Encoder and error propagation to topology reasoning. Therefore, we propose Topology-Inherited Training to inherit better topology knowledge into RoadNetTransformer. Additionally, we collect SD-Maps from open-source map datasets and use this prior information to significantly improve landmark detection and reachability. Extensive experiments on nuScenes dataset demonstrate the superiority of RoadNet Sequence representation and the non-autoregressive approach compared to existing state-of-the-art alternatives. Our code is publicly available at open-source https://github.com/fudan-zvg/RoadNetworkTRansformer.
@@ I Introduction
With the rising prevalence of self-driving cars, a deep knowledge of the road structure is indispensable for autonomous vehicle navigation [1, 2, 3, 4]. Road network [5, 6, 7] extraction is required to estimate highly accurate road landmark locations, centerline curve shapes and road topological connection for self-driving vehicles. However, the ability to understand road network in real-time using onboard sensors is highly challenging.
@@
Unfortunately, existing attempts [17, 18] to extract road network using onboard sensors have not been able to achieve a harmonious integration between Euclidean and non-Euclidean data. STSU [17] divides the road network construction into two stages: center-lines detection and center-line connectivity reasoning—which ignores the cooperation between Euclidean and non-Euclidean domains. TPLR [18] uses a Transformer or Polygon-RNN [19] to combine topology reasoning with center-line localization, but the embedded conflicts between Euclidean data representations and connectivity representation undermine the model performance.
@@
Moreover, the decoupled RoadNet Sequence can be upgrated to accelerate the slow auto-regressive process. Our observation is that the dependency of road network can also be decoupled into a semi-autoregressive format, that retains auto-regressive functionality within local contexts while simultaneously conducting multiple generations in parallel. This semi-autoregressive model is called Semi-Autoregressive RoadNetTransformer (SAR-RNTR). This approach not only accelerates the inference speed by 6 times, but it also significantly boosts the accuracy to a new level based on the better dependency modeling. Going beyond SAR-RNTR, we employ a masked training technique on SAR-RNTR to mimic the remaining auto-regressive dependency through iterative prediction [23]. This gives rise to our Non-Autoregressive RoadNetTransformer (NAR-RNTR) model, which achieves real-time inference speed (47×47$\times$47 × faster) while maintaining the high performance.
To evaluate road network extraction quality, apart from inheriting the former metrics [17] based on lane detection, we instead design a family of metrics directly based on definition of road network – Landmark Precision-Recall and Reachability-Precision-Recall – to evaluate (i) road landmarks localization accuracy, and (ii) path accuracy between any reachable landmarks.
We make the following contributions: (i) we introduce RoadNet Sequence, a lossless, efficient and unified representation of both Euclidean and non-Euclidean information from road network. (ii) We propose a Transformer-based RoadNetTransformer which can decode RoadNet Sequence from multiple onboard cameras. (iii) By decoupling auto-regressive dependency of RoadNet Sequence, our proposed Non-autoregressive RoadNetTransformer accelerate the inference speed to real-time while boost the accuracy with a significant step from the auto-regressive model. (iv) Extensive experiments on nuScenes [24] dataset validate the superiority of RoadNet Sequence representation and RoadNetTransformer over the alternative methods with a considerable margin.
@@
In conclusion, we make the following extension: (i) We examine the advantages of decoupled RoadNet Sequence. (ii) We propose a brand-new training strategy, Topology-Inherited Training, to mitigate error propagation in topology learning, enhancing both Euclidean and non-Euclidean data understanding. (iii) We introduce SDMap-RoadNetTransformer, which realizes the inheritance of topology knowledge across modalities. (iv) We collect SD-Map data from OpenStreetMap [27] and treat SD-Maps as a prompt learning problem without altering the model structure or complex data processing.
@@ II Related work
Vision-based ego-car (BEV) feature learning A rising tendency is conducting self-driving downstream tasks [28, 15, 16, 29, 13, 14, 30, 31, 32, 33] under ego-car coordinate frame. To learn ego-car feature learning from onboard cameras,  [34, 35] project image features to ego-car coordinate based on depth estimation. OFT [28], LSS [13] and FIERY [30] predict depth distribution to generate the intermediate 3D representations. [14, 25, 15, 16] resort neural network like Transformer to learn ego-car feature without depth. In order to learn the topology of the road network under the ego-car coordinate frame, we use the straightforward method of applying LSS [13, 29] to extract ego-car features from the multiple onboard cameras.
Road network extraction Researchers have explored the utilization of DNNs to decode and recover maps from aerial images and GPS trajectories [36, 37]. Moreover, STSU [17] first detect centerline from front-view image with Transformer and then predict the association between centerline with MLP layer, followed with a final merge to estimate road network. Based on STSU, TPLR [18] introduces minimal cycle to eliminate ambiguity in its connectivity representation. Existing methods spend great effort to deal with problem in non-Euclidean domain but ignore the cooperation between Euclidean and non-Euclidean. However, we create a Euclidean-nonEuclidean unified representation.
@@
Non-Autoregressvie generation Non-Autoregressive generation for Neural Machine Translation (NMT) [38] has been proposed as a solution to speed up the one-by-one sequential generation process of autoregressive models. [39, 40, 41] utilized knowledge distillation to help NAT model capture target sequence dependency. [23, 42, 43, 44] proposed iterative models that refine the output from the previous iteration or a noised target in a step-wise manner. [45] kept the auto-regressive approach for global modeling, but introduced the parallel output of a few successive words at each time step. [46, 47, 48, 49] used latent variables as intermediates to reduce the dependency on the target sequence. The current state of non-autoregressive NMT models is far from perfect, with a significant gap in performance compared to their auto-regressive counterparts [50]. But in the case of our RoadNet Sequence, the auto-regressive dependency can be decoupled, leading to both acceleration and improved performance.
@@ III Method
[…]
@@ III-B Coupled RoadNet Sequence
From DAG to Directed Forest A typical DAG always has the ambiguity relationship between vertices and edges ($|\mathcal{V}|\nleftrightarrow|\mathcal{E}|$) so that preserving all edges and vertices leads to redundancy due to the repeated use of vertices when indicating edge direction. In comparison, as shown in Figure 3, the Tree structure has a clear relationship between them ($|\mathcal{V}|\leftrightarrow|\mathcal{E}|$) which avoids this redundancy. Hence, we innovatively transform DAG to a Directed Forest $\mathcal{G}_{f}=(\mathcal{V}_{f},\mathcal{E}_{f})$ (a collection of disconnected directed trees). This transformation has the benefit of enabling us to find a bijection between $\mathcal{E}_{f}$ and $\mathcal{V}_{f}\backslash\{v\in\mathcal{V}_{f}\ |\ id(v)=0\}$ (set of all vertices which is not root, where $id(v)$ means incoming degree of the vertex). The bijection $f:\mathcal{V}\rightarrow\mathcal{E}$ can be $f(v)=(Parent(v),v)$ for any vertices $v\in\mathcal{V}_{f}\backslash\{v\in\mathcal{V}_{f}\ |\ id(v)=0\}$ where $Parent(v)$ is its unique parent and $f^{-1}(e)=v_{1}$ for any $e=(v_{0},v_{1})\in\mathcal{E}$.
@@
Topological sorting of Directed Forest We use Depth-first search (DFS) to obtain the sequential order of vertices $\mathcal{V}_{f}$ in the Directed Forest, i.e., $[v_{1},v_{2},\cdots,v_{n}]$. But the Clone point labeled in the last paragraph will not be traversed. Instead, as shown in the bottom of Figure 3, Clone point will be traversed following its original point.
@@
Non-unique sorting In the application of Depth-First Search (DFS), the sequence in which points under the same parent are searched first can result in non-unique sorting. We investigate two strategies to address this: a random ordering strategy and an ordering based on coordinates, specifically prioritizing points closest to the Bird’s Eye View (BEV) front right.
Sequence construction We then use 6 integers to represent each vertex-edge pair. 6 integers are made up with: (i) two integers for location of vertex, i.e., int$(v_{x})$, int$(v_{y})$; (ii) one integer for category $v_{c}$ of of vertex and we set categories as Ancestor, Lineal, Offshoot, Clone; (iii) one integer for the index of parent (None for root), $v_{d}=$Index$(Parent(v))$ where index is its topological order (the Clone vertex share the same index as origin to indicate that they are identical); (iv) two integers for coefficient of Bezier curve of $e=f(v)$, i.e.int$(e_{px})$, int$(e_{py})$. But, simply discretizing $e_{px}$ and $e_{py}$ can be challenging since the Bezier control points may exceed the Bird’s Eye View (BEV) range, and their values may become negative. As a solution, we discretize $e_{px}$ and $e_{py}$ by applying the int function to $(e_{px}+10)$ and $(e_{py}+10)$, respectively, to avoid negative values.
There exists 4 cases to determine the category $v_{c}$: (i) if $v_{i}$ is the root of the tree, the category $v_{c}$ is set as Ancestor, the index of its parent $v_{d}$ is set as None, its coefficient is ignored. (ii) if $v_{i}$ is the first child of its parent, the category $v_{c}$ is set as Lineal, its $v_{d}$ is ignored for its parent is exact $v_{i-1}$. (iii) if $v_{i}$ is not the first child of its parent, the category $v_{c}$ is set as Offshoot, its $v_{d}$ is its parent’s index. (iv) if $v_{i}$ is the cloned child, the category is set as Clone, its $v_{d}$ is its original child. Integer representation of $v_{c}$ is Ancestor: 0, Lineal: 1, Offshoot: 2, Clone: 3.
Analysis The proposed RoadNet Sequence possess the merits of losslessness, efficiency and interaction. The losslessness of RoadNet Sequence is guaranteed by establishing a bijection between edges and vertices (excluding the root) of trees. Our RoadNet Sequence has a complexity of $\mathcal{O}(|\mathcal{V}_{f}|)=\mathcal{O}(|\mathcal{E}|)$, making it the most efficient. A mixture of Euclidean and non-Euclidean data within a local 6-integer clause facilitates full interaction between the Euclidean and non-Euclidean domains.
@@
Sequence embedding Each vertex-edge pair is represented by 6 integers. To prevent embedding conflicts between the 6 integers, we divide them into separate ranges. As a default, we set the embedding size to $576$, which is sufficient to accommodate all the integer ranges.
@@ III-C Decoupled RoadNet Sequence
Sequence construction As shown in Figure 5, the sequence of a decoupled RoadNet Sequence comprises two components: the vertex sequence and the edge sequence.
@@
So with three parameters $([Index(\text{child node}),e_{mx},e_{my}])$ and the position of this triple, the curve shape and direction of an edge can be determined. Due to the variable nature of the out-degree for each vertex in the DAG, the number of edges extended by each vertex may vary. Consequently, we employ a <Split> token to ascertain the end of the edge subsequence corresponding to a given vertex as the parent node. And at the end of the entire edge sequence, we set the <EOE>(End of Edge Sequence) token to determine the termination.
@@
III-D Auto-Regressive RoadNetTransformer

### v3

@@ Abstract
The extraction of road network is essential for the generation of high-definition maps since it enables the precise localization of road landmarks and their interconnections. However, generating road network poses a significant challenge due to the conflicting underlying combination of Euclidean (e.g., road landmarks location) and non-Euclidean (e.g., road topological connectivity) structures. Existing methods struggle to merge the two types of data domains effectively, but few of them address it properly. Instead, our work establishes a unified representation of both types of data domain by projecting both Euclidean and non-Euclidean data into an integer series called RoadNet Sequence. Further than modeling an auto-regressive sequence-to-sequence Transformer model to understand RoadNet Sequence, we decouple the dependency of RoadNet Sequence into a mixture of auto-regressive and non-autoregressive dependency. Building on this, our proposed non-autoregressive sequence-to-sequence approach leverages non-autoregressive dependencies while fixing the gap towards auto-regressive dependencies, resulting in success in both efficiency and accuracy. We further identify two main bottlenecks in the current RoadNetTransformer on a non-overfitting split of the dataset: poor landmark detection limited by the BEV Encoder and error propagation to topology reasoning. Therefore, we propose Topology-Inherited Training to inherit better topology knowledge into RoadNetTransformer. Additionally, we collect SD-Maps from open-source map datasets and use this prior information to significantly improve landmark detection and reachability. Extensive experiments on the nuScenes dataset demonstrate the superiority of RoadNet Sequence representation and the non-autoregressive approach compared to existing state-of-the-art alternatives. Our code is publicly available at open-source https://github.com/fudan-zvg/RoadNetworkTRansformer.
@@ I Introduction
With the rising prevalence of self-driving cars, a deep knowledge of the road structure is indispensable for autonomous vehicle navigation [1, 2, 3, 4]. Road network [5, 6, 7] extraction is required to estimate highly accurate road landmark locations, centerline curve shapes, and road topological connection for self-driving vehicles. However, the ability to understand road network in real-time using onboard sensors is highly challenging.
@@
Unfortunately, existing attempts [17, 18] to extract road network using onboard sensors have not been able to achieve a harmonious integration between Euclidean and non-Euclidean data. STSU [17] divides the road network construction into two stages: center-lines detection and center-line connectivity reasoning—which ignores the cooperation between Euclidean and non-Euclidean domains. TPLR [18] uses a Transformer or Polygon-RNN [19] to combine topology reasoning with center-line localization, but the embedded conflicts between Euclidean data representations and connectivity representations undermine the model performance.
@@
Moreover, the decoupled RoadNet Sequence can be upgrated to accelerate the slow auto-regressive process. Our observation is that the dependency of road network can also be decoupled into a semi-autoregressive format, that retains auto-regressive functionality within local contexts while simultaneously conducting multiple generations in parallel. This semi-autoregressive model is called Semi-Autoregressive RoadNetTransformer (SAR-RNTR). This approach not only accelerates the inference speed by 6 times, but it also significantly boosts the accuracy to a new level based on the better dependency modeling. Going beyond SAR-RNTR, we employ a masked training technique on SAR-RNTR to mimic the remaining auto-regressive dependency through iterative prediction [23]. This gives rise to our Non-Autoregressive RoadNetTransformer (NAR-RNTR) model, which achieves real-time inference speed ($47\times$ faster) while maintaining the high performance.
To evaluate road network extraction quality, apart from inheriting the former metrics [17] based on lane detection, we instead design a family of metrics directly based on the definition of the road network – Landmark Precision-Recall and Reachability-Precision-Recall – to evaluate (i) road landmarks localization accuracy, and (ii) path accuracy between any reachable landmarks.
We make the following contributions: (i) we introduce RoadNet Sequence, a lossless, efficient and unified representation of both Euclidean and non-Euclidean information from road network. (ii) We propose a Transformer-based RoadNetTransformer which can decode RoadNet Sequence from multiple onboard cameras. (iii) By decoupling auto-regressive dependency of RoadNet Sequence, our proposed Non-autoregressive RoadNetTransformer accelerates the inference speed to real-time while boosting the accuracy with a significant step from the auto-regressive model. (iv) Extensive experiments on nuScenes [24] dataset validate the superiority of RoadNet Sequence representation and RoadNetTransformer over the alternative methods with a considerable margin.
@@
In conclusion, we make the following extension: (i) We examine the advantages of decoupled RoadNet Sequence; (ii) We propose a brand-new training strategy, Topology-Inherited Training, to mitigate error propagation in topology learning, enhancing both Euclidean and non-Euclidean data understanding; (iii) We introduce SDMap-RoadNetTransformer, which realizes the inheritance of topology knowledge across modalities; (iv) We collect SD-Map data from OpenStreetMap [27] and treat SD-Maps as a prompt learning problem without altering the model structure or complex data processing.
@@ II Related work
Vision-based ego-car (BEV) feature learning. A rising tendency is conducting self-driving downstream tasks [28, 15, 16, 29, 13, 14, 30, 31, 32, 33] under ego-car coordinate frame. To learn ego-car feature learning from onboard cameras,  [34, 35] project image features to ego-car coordinate based on depth estimation. OFT [28], LSS [13] and FIERY [30] predict depth distribution to generate the intermediate 3D representations. [14, 25, 15, 16] resort neural network like Transformer to learn ego-car feature without depth. In order to learn the topology of the road network under the ego-car coordinate frame, we use the straightforward method of applying LSS [13, 29] to extract ego-car features from the multiple onboard cameras.
Road network extraction. Researchers have explored the utilization of DNNs to decode and recover maps from aerial images and GPS trajectories [36, 37]. Moreover, STSU [17] first detects centerline from front-view image with Transformer and then predicts the association between centerline with MLP layer, followed by a final merge to estimate road network. Based on STSU, TPLR [18] introduces minimal cycle to eliminate ambiguity in its connectivity representation. Existing methods spend great effort to deal with the problem in the non-Euclidean domain but ignore the cooperation between Euclidean and non-Euclidean. Previous approaches, such as STSU [17] and TPLR [18], focus on solving the problem within the non-Euclidean domain, typically handling road network connectivity through centerline detection or minimal cycles. While these methods tackle topological reasoning and road connectivity, they often neglect the interdependencies between Euclidean and non-Euclidean information. For example, STSU detects centerlines using Transformer-based models and then predicts the association between centerlines through an MLP, but it does not effectively merge road geometry with topological structure. TPLR addresses connectivity ambiguities by introducing minimal cycles, but it does not fully integrate Euclidean data with non-Euclidean topological data. In contrast, our work proposes a unified representation that integrates both Euclidean and non-Euclidean road network information into a single framework, overcoming the limitations of previous methods. Instead of treating these domains separately, as done by STSU [17] and TPLR [18], we introduce a new method that projects both Euclidean and non-Euclidean data into an integer series called RoadNet Sequence. By doing so, we create a unified, lossless, and interactive representation that allows for more accurate modeling of road networks. Further, we introduce a sequence-to-sequence Transformer model that decouples the dependency of RoadNet Sequence into a mixture of auto-regressive and non-autoregressive dependencies. This innovative approach allows our model to leverage the benefits of both dependencies, ensuring high efficiency and accuracy in predicting road network properties. Our proposed non-autoregressive approach significantly improves the efficiency of training and inference while maintaining the model’s predictive accuracy.
@@
Non-Autoregressvie generation. Non-Autoregressive generation for Neural Machine Translation (NMT) [38] has been proposed as a solution to speed up the one-by-one sequential generation process of autoregressive models. [39, 40, 41] utilized knowledge distillation to help the NAT model capture target sequence dependency. [23, 42, 43, 44] proposed iterative models that refine the output from the previous iteration or a noised target in a step-wise manner. [45] kept the auto-regressive approach for global modeling, but introduced the parallel output of a few successive words at each time step. [46, 47, 48, 49] used latent variables as intermediates to reduce the dependency on the target sequence. The current state of non-autoregressive NMT models is far from perfect, with a significant gap in performance compared to their auto-regressive counterparts [50]. But in the case of our RoadNet Sequence, the auto-regressive dependency can be decoupled, leading to both acceleration and improved performance.
@@ III Method
Figure 3: Top illustrates the transformation process from a Directed Acyclic Graph to a Directed Forest for RoadNet Sequence. The red boxes enclose all merge-points that have non-unique parents. We replicate theirs parents (except the first parent) to their children (cyan nodes) and delete the corresponding edges. Middle demonstrates a bijection between the edges and vertices ($\dagger$}: excluding the Ancestors). Bottom lists the topological sorting result of vertices and presents six integers for each vertex.
@@ III-B Coupled RoadNet Sequence
[…]
@@
Topological sorting of Directed Forest. We use Depth-first search (DFS) to obtain the sequential order of vertices $\mathcal{V}_{f}$ in the Directed Forest, i.e., $v_{1},v_{2},\cdots,v_{n}]$. But the Clone point labeled in the last paragraph will not be traversed. Instead, as shown in the bottom of Figure 3, Clone point will be traversed following its original point.
@@
Non-unique sorting. In the application of Depth-First Search (DFS), the sequence in which points under the same parent are searched first can result in non-unique sorting. We investigate two strategies to address this: a random ordering strategy and an ordering based on coordinates, specifically prioritizing points closest to the Bird’s Eye View (BEV) front right.
[…]
There exist 4 cases to determine the category $v_{c}$: (i) if $v_{i}$ is the root of the tree, the category $v_{c}$ is set as Ancestor, the index of its parent $v_{d}$ is set as None, its coefficient is ignored. (ii) if $v_{i}$ is the first child of its parent, the category $v_{c}$ is set as Lineal, its $v_{d}$ is ignored for its parent is exact $v_{i-1}$. (iii) if $v_{i}$ is not the first child of its parent, the category $v_{c}$ is set as Offshoot, its $v_{d}$ is its parent’s index. (iv) if $v_{i}$ is the cloned child, the category is set as Clone, its $v_{d}$ is its original child. Integer representation of $v_{c}$ is Ancestor: 0, Lineal: 1, Offshoot: 2, Clone: 3.
Details of sequence construction. The discretization of $v_{x},v_{y}$ is simply truncating the integer part. Integer representation of $v_{c}$ is Ancestor: 0, Lineal: 1, Offshoot: 2, Clone: 3. Discretizing $e_{px}$ and $e_{py}$ can be challenging since the Bezier control points may exceed the Bird’s Eye View (BEV) range, and their values may become negative. As a solution, we discretize $e_{px}$ and $e_{py}$ by applying the int function to $(e_{px}+10)$ and $(e_{py}+10)$, respectively, to avoid negative values.
Analysis. The proposed RoadNet Sequence possess the merits of losslessness, efficiency and interaction. The losslessness of RoadNet Sequence is guaranteed by establishing a bijection between edges and vertices (excluding the root) of trees. Our RoadNet Sequence has a complexity of $\mathcal{O}(|\mathcal{V}_{f}|)=\mathcal{O}(|\mathcal{E}|)$, making it the most efficient. A mixture of Euclidean and non-Euclidean data within a local 6-integer clause facilitates full interaction between the Euclidean and non-Euclidean domains.
The transformation from a Directed Acyclic Graph (DAG) to a Directed Forest shares similarities with the lookahead method in Transformers, especially in capturing sequential dependencies and enhancing future predictions. However, the key difference lies in the methodology: while the lookahead method in Transformers uses dynamic attention to model dependencies, our DAG-to-Directed Forest transformation simplifies the graph structure to reduce redundancy, making the prediction process more efficient. Additionally, while Transformers rely on embeddings and attention, our method uses a predefined structural transformation tailored for road network sequence prediction.
@@
Sequence embedding. Each vertex-edge pair is represented by 6 integers. To prevent embedding conflicts between the 6 integers, we divide them into separate ranges. As a default, we set the embedding size to $576$, which is sufficient to accommodate all the integer ranges.
@@ III-C Decoupled RoadNet Sequence
Sequence construction. As shown in Figure 5, the sequence of a decoupled RoadNet Sequence comprises two components: the vertex sequence and the edge sequence.
@@
[…]
@@
III-D Input and Target Sequence Construction
Sequence embedding. Each vertex-edge pair is represented by 6 integers. To prevent embedding conflicts between the 6 integers, we divide them into separate ranges which is shown in Table I. As a default, we set the embedding size to $576$, which is sufficient to accommodate all the integer ranges.
TABLE I: Embedding range of different integers.
Item Range
$v_{x},v_{y}$ $0\sim 199$
$v_{c}$ $200\sim 249$
$v_{d}$ $250\sim 349$
$e_{px},e_{py}$ $350\sim 569$
noise category $570$
EOS $571$
Start 57572
n/a $573$
Synthetic noise objects technique. The input sequence of RoadNet Sequence starts with a start token and the target sequence ends with an EOS token. The EOS token makes the model know where the sequence terminates, but the experiments have shown that it tends to cause the model to stop predicting early without getting the complete sequence. Inspired by [22], we use a similar sequence augmentation technique to alleviate the problem called the synthetic noise objects technique [22]. The technique composes of sequence augmentation and sequence noise padding. The sequence augmentation adds noise to the position of landmarks and the coefficient of centerlines in input sequence. Whereas, sequence noise padding is a padding technique. For input sequences, we generate synthetic noise vertices and append them at the end of the real vertices sequence. Each noise vertex includes random locations($v_{x}$, $v_{y}$), category($v_{c}$), index of parent($v_{d}$) and Bezier curve coefficient($e_{px}$, $e_{py}$). As for the target sequence, the EOS token is added to the end of the real vertices sequence. We set the target category($v_{c}$) of each noise vertex to a specific noise class(different from any of the ground-truth labels), and the remaining components($v_{x}$, $v_{y}$, $v_{d}$, $e_{px}$, $e_{py}$) of the noise vertex to the ”n/a” class, whose loss is not calculated in the back-propagation.
However, we only use sequence noise padding as sequence augmentation has been shown to cause a decrease in performance. The introduced modifications of the synthetic noise objects technique are illustrated in Figure 6.
The padding rules of Semi-Autoregressive RoadNet Sequence are much the same as auto-regressive RoadNet Sequence. As mentioned in the main submission, we pad the 2-dimensional Semi-Autoregressive RoadNet Sequence to $y_{1,1},y_{1,2},\cdots,y_{1,L}],\cdots,[y_{M,1},y_{M,2},\cdots,y_{M,L}]]$, where $L$ is the maximum length of each sub-sequence and $M$ is the number of sub-sequences. The valid sub-sequences begin with a key-point. For each valid sub-sequence, we follow the same padding rules of RoadNet Sequence, except there isn’t a start token in an input sub-sequence because of the Key-point Prompt. We set the other sub-sequences to the ”n/a” class making the loss of these sub-sequences without a key-point not calculated.
III-E Auto-Regressive RoadNetTransformer